OpenStudy (anonymous):

a rectangular table is two times as long as it is wide. if the area is 50ft^2, find the length and the width of the table.

OpenStudy (jlastino):

Let x= width 2x= length Area of rectangle = length(width) 50 =x(2x) 2x^2= 50 x^2=25 x=5 2x= 10 Length = 10ft Width = 5ft